Scot and Vidia, married taxpayers, earn $254,400 in taxable income and $5,000 in interest from an investment in City of Tampa bonds. (Use the U.S. tax rate schedule for married filing jointly).

Schedule Y-1-Married Filing Jointly or Qualifying Widow(er)

If taxable income is over: But not over: The tax is:
$          0 $ 19,750 10% of taxable income
$ 19,750 $ 80,250 $1,975 plus 12% of the excess over $19,750
$ 80,250 $171,050 $9,235 plus 22% of the excess over $80,250
$171,050 $326,600 $29,211 plus 24% of the excess over $171,050
$326,600 $414,700 $66,543 plus 32% of the excess over $326,600
$414,700 $622,050 $94,735 plus 35% of the excess over $414,700
$622,050 $167,307.50 plus 37% of the excess over $622,050

Required:

  

  1. If Scot and Vidia earn an additional $86,000 of taxable income, what is their marginal tax rate on this income?
  2. What is their marginal tax rate if, instead, they report an additional $86,000 in deductions?
Answer #1
Tax on $ 254,400     49,215 {29,211 + ( ( 254,400 - 171,050 ) x 24% ) }
a.) Additional Income     86,000
Total taxable income $ 340,400 =254400+86000
Tax on $ 340,400     70,959 =66543+((340400-326600)*32%)
Increase in Tax     21,744 (70,959 - 49,215 )
Marginal Tax rate 25.28% (21,744 / 86,000 )
b.) Additional Deduction     86,000
Total taxable income $ 168,400 =254400-86000
Tax on $ 168,400     28,628 {9,235 + ( ( 168,400 - 80,250 ) x 22% ) }
Decrease in Tax     20,587 (49,215 - 28,628 )
Marginal Tax rate 23.94% (20,587 / 86,000 )